# GridPE: Unifying Positional Encoding in Transformers with a Grid Cell-Inspired Framework

### v1

@@ Abstract
Understanding spatial location and relationships is a fundamental capability for modern artificial intelligence systems. Insights from human spatial cognition provide valuable guidance in this domain. Recent neuroscientific discoveries have highlighted the role of grid cells as a fundamental neural component for spatial representation, including distance computation, path integration, and scale discernment. In this paper, we introduce a novel positional encoding scheme inspired by Fourier analysis and the latest findings in computational neuroscience regarding grid cells. Assuming that grid cells encode spatial position through a summation of Fourier basis functions, we demonstrate the translational invariance of the grid representation during inner product calculations. Additionally, we derive an optimal grid scale ratio for multi-dimensional Euclidean spaces based on principles of biological efficiency. Utilizing these computational principles, we have developed a Grid-cell inspired Positional Encoding technique, termed GridPE, for encoding locations within high-dimensional spaces. We integrated GridPE into the Pyramid Vision Transformer architecture. Our theoretical analysis shows that GridPE provides a unifying framework for positional encoding in arbitrary high-dimensional spaces. Experimental results demonstrate that GridPE significantly enhances the performance of transformers, underscoring the importance of incorporating neuroscientific insights into the design of artificial intelligence systems.
1 Introduction
@@
Understanding how the mammalian brain represents spatial characteristics and relationships remains a critical question in cognitive science and neuroscience. Specifically, the focus is on identifying the neural bases and their roles in self-localization, navigation, and path planning. Scholars have conducted extensive research to address this question. [1] designed an animal experiment where rats performed a wayfinding task in a radial maze, introducing the concept of the cognitive map to describe the internal environmental representation that supports flexible spatial behavior. Subsequent discoveries identified the neural cellular basis of the cognitive map in the mammalian brain, a line of research that earned the Nobel Prize in 2014. Place cells activate when a mammal enters an area near a specific spatial location [2, 3], and are believed to distinguish and represent different places. Grid cells in the medial entorhinal cortex (MEC), on the other hand, exhibit periodic, hexagonal, and grid-like firing fields [4]. These cells are thought to define distance metrics in Euclidean space, support multi-scale self-localization, and facilitate path integration and vector-based navigation [5, 6, 7]. In other words, grid cells represent spatial relationships independently of the agent’s current visual sensory input and head direction (referred to as "allocentric" representation). Due to their close association with spatial representation, the principles of grid cell firing have been widely adopted by artificial intelligence researchers beyond neuroscience, including in robotics navigation system design [8], reinforcement learning agent trajectory planning [9], and positional encoding for geospatial data [10, 11].
@@
There are at least two approaches to incorporating grid cells into AI systems. For bottom-up approach, agents can learn from trajectory experience (i.e., model training in the context of machine learning), and grid-like firing patterns can often emerge from certain neuron activations in the architecture [7, 12]. While these agent architectures validate the existence of grid cells, they struggle to directly encode spatial positions from knowledge of the model. For the top-down methods, researchers intentionally design computational models to simulate grid cell firing and integrate them into the agent for navigation. Two commonly adopted techniques are the continuous attractor network (CAN) and the velocity-controlled oscillator (VCO). The CAN model constructs a recurrent neural network that takes velocity input to compute a steady state from this neural dynamic system, simulating grid cells [13]. However, CAN presumes symmetric connections and regular cellular arrangement in the network, which rarely hold true in authentic neural anatomical structure [14]. VCO models assume that grid cell firing patterns arise from the interference of oscillations with distinct frequencies [15]. A grid cell’s activation pattern can be the summation of three planar waves with respective frequency direction vectors, each separated by 120°[16]. VCO models encode location with high simplicity and are theoretically supported by neuroscience evidence such as band cells [17] and phase precession phenomena in bats [18, 19].
@@
Inspired by the spatial encoding properties of grid cells, we designed a positional encoding scheme named GridPE for representing multi-dimensional spatial positions in self-attention architectures. To compare GridPE with classical positional encoding techniques, we conducted experiments using the Pyramid Vision Transformer (PVT) architecture [20] on the ImageNet-1000 image classification task [21]. The results demonstrate its effectiveness in enhancing the self-attention mechanism by providing both absolute and relative positional information, highlighting the potential of integrating neuroscience findings into AI system design.
2 Preliminaries
We model the grid cell on the theoretical basis of VCO theory [15, 22]. Under this assumption, presynaptic cells of the grid cell (sometimes referred to as "simple cells" [14]) perform oscillations similar to two-dimensional planar waves, forming parallel band "strips" along specific directions. The oscillation frequency of the presynaptic dendrites is sensitive to the agent’s current velocity, which is why the theory is named "velocity-controlled." The frequency can be expressed as the result of modulating the somatic oscillation base frequency, with the modulation magnitude depending on the velocity vector and the presynaptic preferred direction (or wave vector of the planar wave) [9].
@@ 2 Preliminaries
where $\omega_{s}$ is the somatic base frequency, usually experimentally measured as theta wave frequency (4-11 Hz) [9]. $\boldsymbol{v},\boldsymbol{k}\in\mathbb{R}^{2}$ are the velocity vector and wave vector in the two-dimensional space, the latter indicating preferred direction and wave length. $\beta$ is a position factor deciding sensitivity to velocity. The velocity-modulated dendrite frequency $\omega_{d}$ is the derivative of oscillation phase $\phi$. Here, we do not differentiate between angular velocity and frequency due to their constant ratio of $2\pi$. When we integrate the modulated frequency over the time $t$, we obtain the oscillation phase:
@@
where $\boldsymbol{x}$ is the displacement vector from time 00 to time $t$. Clearly, at a given time, locations with the same projection onto the wave vector correspond to the same oscillation phase, thus form a parallel activation bands across the environments when we calculate amplitude. If we overlap multiple such presynaptic firing fields with weights, we can obtain grid cell hexagonal firing patterns. The oscillation is expressed as Fourier basis function, and grid cell activation at any location can be expressed as [23]:
$g$ $=\sum_{i=0}^{n}c_{i}\,e^{j\omega_{s}t}\cdot e^{j\beta\boldsymbol{k}_{i}^{T}\boldsymbol{x}}$ (3)
$=\sum_{i=0}^{n}c_{i}(t)e^{j\boldsymbol{k}_{i}^{T}\boldsymbol{x}}$
Imagery unit is denoted as $j$, and the number of presynampic cells is denoted as $n$. We omit the $\beta$ which can be integrated into wave vector $k$. For the second equality, the time-dependent function $c_{i}(t)$ combine the information of synaptic weight and somatic base oscillation, while the factor of Fourier basis only depend on location $x$. That means the temporal fluctuation of the activation is only driven by the intrinsic theta waves evolving as time progresses, which is independent from spatial representation. Therefore, we can fix to a certain time point and discuss the grid cell representation across the space from a static viewpoint, for which we omit the time variable $t$ in the following derivation. In previous research, three Fourier bases with their directional preferences forming angles of 120° between each other are commonly used [23, 16], as illustrated in Figure 1.
@@
3 Derivation of Grid Cell Positional Encoding
3.1 Inner Product Translational Invariance
@@
Note this function is also a translational invariant kernel for $\boldsymbol{x}$ and $\boldsymbol{y}$. Simultaneuously, in its form it scale to the inverse Fourier transformation of the probability distribution of wave vectors. According to Bochner theorem [24], we know this kernel, as the inverse Fourier transformation of a non-nagative measurement, is positive definite. Therefore, we can apply random Fourier feature (RFF) [25] technique on this shift-invariant kernel, which demands positive definite kernel as the prerequisite. We denote the $\kappa(\boldsymbol{x},\boldsymbol{y})$ as this kernel and we can know:
@@
$=\frac{1}{\sqrt{D}}\left(\begin{array}[]{c}e^{j\boldsymbol{k}_{1}^{T}\boldsymbol{x}}\\ e^{j\boldsymbol{k}_{2}^{T}\boldsymbol{x}}\\ \vdots\\ e^{j\boldsymbol{k}_{D}^{T}\boldsymbol{x}}\end{array}\right)^{T}\begin{bmatrix}|c_{1}|^{2}&&&\\ &|c_{2}|^{2}&&\\ &&\ddots&\\ &&&|c_{D}|^{2}\end{bmatrix}\left(\begin{array}[]{c}e^{-j\boldsymbol{k}_{1}^{T}\boldsymbol{y}}\\ e^{-j\boldsymbol{k}_{2}^{T}\boldsymbol{y}}\\ \vdots\\ e^{-j\boldsymbol{k}_{D}^{T}\boldsymbol{y}}\end{array}\right)\frac{1}{\sqrt{D}}$
@@
We propose that the brain represents Euclidean space using multi-dimensional Fourier analysis, akin to the auditory system’s information processing. By integrating oscillations of various spatial frequencies, the grid cell encoding system can efficiently represent spatial information with high resolution. Furthermore, the Random Fourier Features (RFF) approach imposes no constraints on the dimensionality of locations $\boldsymbol{x}$ and $\boldsymbol{y}$. This suggests that the Fourier representation of grid cells can encode location information in arbitrarily high-dimensional spaces.
3.2 Optimal Ratio of Grid Scale under Economy Principle
In the derivation above, we assumed the wave vector (i.e., the preferred direction) follows a probability distribution, such as the uniform distribution often used in RFF. However, neuroscience evidence shows that the medial entorhinal cortex consists of several modules. Within the same module, the direction and period scale of the cell grids remain consistent, but these discretized scales shrink along the ventral-dorsal axis [26], representing a transition from coarse, low-frequency to granular, high-frequency spatial information. The periodic ratio of the adjacent scale remains around 1.5, as estimated by both animal measurements and computational simulations [26, 7]. This hierarchical spatial representation demands significantly fewer grid cells while maintaining fine spatial resolution, or say the economy principle [27].
The challenge now is to determine the optimal ratio of the grid scale, which is crucial for constructing a positional encoding system. Recall that in the original positional encoding of the Transformer architecture [28], the encoding spatial frequency is calculated as $\omega_{k}=\frac{1}{10000^{2k/d}}$, forming a geometric series of wavelengths. The rationale behind this choice remains unclear, and we anticipate that computational theories of grid cells can provide insights into frequency determination.
@@ 3.2 Optimal Ratio of Grid Scale under Economy Principle
Here we provide a generalized derivation based on the proof in [27]. We aim to represent p-dimensional Euclidean space with the least number of grid cells, reflecting the principle of economy. Assume there are $m$ scales (or layers) of the grid cells. The period of the $i$-th module layer ($i=1,2,\ldots,m$), or the distance between the adjacent grid vertices, is denoted as $\lambda_{i}$. The period increase as $i$ becomes larger. The scale ratio of the adjacent layers is $r=\lambda_{i+1}/\lambda_{i}>1$. A grid field is a hyperspherical area centered around a grid vertex in the space, with its diameter denoted as $l_{i}$. For 2D space, a grid field is a circular-plate area activating the grid cell. Clearly, to cover the whole space without omissions, more than one grid cell with different phases are required. The larger the ratio of $\lambda_{i}$ to $l_{i}$, the more grid cells are needed. If there must be at least $d$ cells responding to a spatial location, the total number of grid cells across all layers can be expressed as $N$:
@@
3.3 Proposal of Grid-cell Positional Encoding (GridPE)
After demonstrating the capability of representing relative movement in space through inner product calculations and determining the optimal ratio of encoding periods, we introduce a novel positional encoding scheme for attention-based architectures. Inspired by neuroscientific principles of grid cells, our method extends the foundational ideas presented by  [29]. While Su’s model combines absolute and relative positional encodings within a fixed framework, our schema advances this by incorporating grid cell-inspired encodings that inherently support extrapolation to higher dimensions and maintain translational invariance. Notably, in one-dimensional space, where GridPE can be expressed using a single Fourier basis, our encoding scheme is equivalent to rotational positional encoding. This demonstrates the flexibility of our approach and its alignment with established encoding strategies, further enhancing its applicability across various architectural contexts. According to Su’s model, to accommodate both absolute and relative positional encoding, the inner product of the query $\boldsymbol{q}$ and key $\boldsymbol{k}$, which represents their similarity, can be modified as follows:
@@ 3.3 Proposal of Grid-cell Positional Encoding (GridPE)
$\boldsymbol{W}_{m}=\left[\sum_{i}c_{i}e^{j\boldsymbol{\omega}_{i,1}^{T}m},\sum_{i}c_{i}e^{j\boldsymbol{\omega}_{i,2}^{T}m},\dots,\sum_{i}c_{i}e^{j\boldsymbol{\omega}_{i,d}^{T}m}\right]^{T}$ (12)
@@
$\boldsymbol{R}_{m}=\begin{bmatrix}\Re\left(e^{j\boldsymbol{\omega}_{1,1}^{T}m}\right)&-\Im\left(e^{j\boldsymbol{\omega}_{1,1}^{T}m}\right)&0&\dots&0&0\\ \Im\left(e^{j\boldsymbol{\omega}_{1,1}^{T}m}\right)&\Re\left(e^{j\boldsymbol{\omega}_{1,1}^{T}m}\right)&0&\dots&0&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&0&\dots&\Re\left(e^{j\boldsymbol{\omega}_{d/2,1}^{T}m}\right)&-\Im\left(e^{j\boldsymbol{\omega}_{d/2,1}^{T}m}\right)\\ 0&0&0&\dots&\Im\left(e^{j\boldsymbol{\omega}_{d/2,1}^{T}m}\right)&\Re\left(e^{j\boldsymbol{\omega}_{d/2,1}^{T}m}\right)\\ \end{bmatrix}$ (13)
@@
We can directly use the real part of the inner product between two sets of "complex" position encoding as the encoding product.
@@
4 Experiments
4.1 Property of Distance Decay
We firstly display that our positional encoding approach exhibits the property of distance decay in any high-dimensional space. This implies that the association between two positions diminishes as the distance between them increases. We select a square area in two-dimensional space, calculating the positional encoding vector within the area as Equation (12), and compute the inner products between the central "core" point and uniformly sampled points. Given that we have proven the translational invariance of the inner product, our findings remain valid when any other point is chosen as the core. Figure 3 presents the inner product surface plot and a profile plot with $x$ fixed at $0.5$.

### v2

@@ Abstract
Understanding spatial location and relationships is a fundamental capability for modern artificial intelligence systems. Insights from human spatial cognition provide valuable guidance in this domain. Neuroscientific discoveries have highlighted the role of grid cells as a fundamental neural component for spatial representation, including distance computation, path integration, and scale discernment. In this paper, we introduce a novel positional encoding scheme inspired by Fourier analysis and the latest findings in computational neuroscience regarding grid cells. Assuming that grid cells encode spatial position through a summation of Fourier basis functions, we demonstrate the translational invariance of the grid representation during inner product calculations. Additionally, we derive an optimal grid scale ratio for multi-dimensional Euclidean spaces based on principles of biological efficiency. Utilizing these computational principles, we have developed a Grid-cell inspired Positional Encoding technique, termed $GridPE$, for encoding locations within high-dimensional spaces. We integrated GridPE into the Pyramid Vision Transformer architecture. Our theoretical analysis shows that GridPE provides a unifying framework for positional encoding in arbitrary high-dimensional spaces. Experimental results demonstrate that GridPE significantly enhances the performance of transformers, underscoring the importance of incorporating neuroscientific insights into the design of artificial intelligence systems. Our code can be accessed via the following anonymous link: https://anonymous.4open.science/r/gridpe-vision-DCF0/
Introduction
@@
Understanding how the mammalian brain represents spatial characteristics and relationships remains a critical question in cognitive science and neuroscience. Specifically, the focus is on identifying the neural bases and their roles in self-localization, navigation, and path planning. Scholars have conducted extensive research to address this question. (Tolman 1948) designed an animal experiment where rats performed a wayfinding task in a radial maze, introducing the concept of the cognitive map to describe the internal environmental representation that supports flexible spatial behavior. Subsequent discoveries identified the neural cellular basis of the cognitive map in the mammalian brain, a line of research that earned the Nobel Prize in 2014. Place cells activate when a mammal enters an area near a specific spatial location (O’keefe and Nadel 1978; Solstad, Moser, and Einevoll 2006), and are believed to distinguish and represent different places. Grid cells in the medial entorhinal cortex (MEC), on the other hand, exhibit periodic, hexagonal, and grid-like firing fields (Hafting et al. 2005). These cells are thought to define distance metrics in Euclidean space, support multi-scale self-localization, and facilitate path integration and vector-based navigation (Gorchetchnikov and Grossberg 2007; Moser, Kropff, and Moser 2008; Banino et al. 2018). In other words, grid cells represent spatial relationships independently of the agent’s current visual sensory input and head direction (referred to as ”allocentric” representation). Due to their close association with spatial representation, the principles of grid cell firing have been widely adopted by artificial intelligence researchers beyond neuroscience, including in robotics navigation system design (Yuan et al. 2015), reinforcement learning agent trajectory planning (Yu, Behrens, and Burgess 2020), and positional encoding for geospatial data (Mai et al. 2020, 2023).
@@
There are at least two approaches to incorporating grid cells into AI systems. For bottom-up approach, agents can learn from trajectory experience (i.e., model training in the context of machine learning), and grid-like firing patterns can often emerge from certain neuron activations in the architecture (Banino et al. 2018; Whittington et al. 2022). While these agent architectures validate the existence of grid cells, they struggle to directly encode spatial positions from knowledge of the model. For the top-down methods, researchers intentionally design computational models to simulate grid cell firing and integrate them into the agent for navigation. Two commonly adopted techniques are the continuous attractor network (CAN) and the velocity-controlled oscillator (VCO). The CAN model constructs a recurrent neural network that takes velocity input to compute a steady state from this neural dynamic system, simulating grid cells (Burak and Fiete 2009). However, CAN presumes symmetric connections and regular cellular arrangement in the network, which rarely hold true in authentic neural anatomical structure (Anselmi, Murray, and Franceschiello 2020). VCO models assume that grid cell firing patterns arise from the interference of oscillations with distinct frequencies (Burgess, Barry, and O’keefe 2007). A grid cell’s activation pattern can be the summation of three planar waves with respective frequency direction vectors, each separated by 120°(Blair, Welday, and Zhang 2007). VCO models encode location with high simplicity and are theoretically supported by neuroscience evidence such as band cells (Krupic, Burgess, and O’Keefe 2012) and phase precession phenomena in bats (Eliav et al. 2018; Yartsev, Witter, and Ulanovsky 2011).
@@
Inspired by the spatial encoding properties of grid cells, we designed a positional encoding scheme named GridPE for representing multi-dimensional spatial positions in self-attention architectures. To compare GridPE with classical positional encoding techniques, we conducted experiments using classic architectures such as Vision Transformer (ViT) and Pyramid Vision Transformer (PVT) architecture (Wang et al. 2021) on the ImageNet-100 image classification task (Deng et al. 2009). The results demonstrate its effectiveness in enhancing the self-attention mechanism by providing both absolute and relative positional information, highlighting the potential of integrating neuroscience findings into AI system design.
Preliminaries
We model the grid cell on the theoretical basis of VCO theory (Burgess, Barry, and O’keefe 2007; Hasselmo 2008). Under this assumption, presynaptic cells of the grid cell (sometimes referred to as ”simple cells” (Anselmi, Murray, and Franceschiello 2020)) perform oscillations similar to two-dimensional planar waves, forming parallel band ”strips” along specific directions. The oscillation frequency of the presynaptic dendrites is sensitive to the agent’s current velocity, which is why the theory is named ”velocity-controlled.” The frequency can be expressed as the result of modulating the somatic oscillation base frequency, with the modulation magnitude depending on the velocity vector and the presynaptic preferred direction (or wave vector of the planar wave) (Yu, Behrens, and Burgess 2020).
@@
where $\omega_{s}$ is the somatic base frequency, usually experimentally measured as theta wave frequency (4-11 Hz) (Yu, Behrens, and Burgess 2020). $\boldsymbol{v},\boldsymbol{k}\in\mathbb{R}^{2}$ are the velocity vector and wave vector in the two-dimensional space, the latter indicating preferred direction and wave length. $\beta$ is a position factor deciding sensitivity to velocity. The velocity-modulated dendrite frequency $\omega_{d}$ is the derivative of oscillation phase $\phi$. Here, we do not differentiate between angular velocity and frequency due to their constant ratio of $2\pi$. When we integrate the modulated frequency over the time $t$, we obtain the oscillation phase:
@@
Where $\boldsymbol{x}$ is the displacement vector from time 00 to time $t$. Clearly, at a given time, locations with the same projection onto the wave vector correspond to the same oscillation phase, thus forming parallel activation bands across the environment when we calculate the amplitude. If we overlap multiple such presynaptic firing fields with weights, we can obtain grid cell hexagonal firing patterns. The oscillation is expressed as a Fourier basis function, and grid cell activation at any location can be expressed as (Dang et al. 2021):
$g$ $=\sum_{i=1}^{n}c_{i}\cos(\varphi_{i}(t))$ (3)
$=\sum_{i=1}^{n}\frac{c_{i}}{2}\left(e^{j\varphi_{i}(t)}+e^{-j\varphi_{i}(t)}\right)$
Here, for the Fourier basis transformation, we consider only the positive frequency components and ignore the negative frequency components. In the context of grid cells, the positive frequency components are sufficient to describe the spatial periodicity and the phase relationships that give rise to the hexagonal grid pattern. The negative frequency components do not provide additional unique information about the grid cell firing patterns. By focusing on the positive frequency components, we simplify the mathematical representation without losing any essential information about the grid cell activation.
$g$ $=\sum_{i=1}^{n}\frac{c_{i}}{2}e^{j\varphi_{i}(t)}$ (4)
$=\sum_{i=1}^{n}\frac{c_{i}}{2}e^{j(\omega_{s}t+\beta\boldsymbol{k}_{i}^{T}\boldsymbol{x})}$
$=\sum_{i=1}^{n}\frac{c_{i}}{2}e^{j\omega_{s}t}\cdot e^{j\beta\boldsymbol{k}_{i}^{T}\boldsymbol{x}}$
$=\sum_{i=1}^{n}\tilde{c}_{i}(t)e^{j\boldsymbol{k}_{i}^{T}\boldsymbol{x}}$
The imagery unit is denoted as $j$, and the number of presynaptic cells is denoted as $n$. We omit the parameter $\beta$, which can be integrated into the wave vector $k$. The time-dependent function $\tilde{c}_{i}(t)$ combines the information of synaptic weight and somatic base oscillation, while the factor of the Fourier basis depends only on the location $x$. This implies that the temporal fluctuation of the activation is driven solely by the intrinsic theta waves evolving over time, which are independent of spatial representation. Consequently, we can fix a certain time point and analyze the grid cell representation across space from a static viewpoint, omitting the time variable $t$ in the subsequent derivation.
Previous research commonly utilizes three Fourier bases with directional preferences forming 120° angles between each other (Dang et al. 2021; Blair, Welday, and Zhang 2007), as illustrated in Figure 1. This arrangement of grid cells forming 120° angles results in a hexagonal lattice structure. Such a hexagonal arrangement is advantageous because it allows for the most efficient and uniform coverage of two-dimensional space, minimizing overlap and maximizing the distinctness of each cell’s spatial representation. This geometric efficiency is crucial for robust spatial navigation and representation, providing a highly regular and repeating pattern that facilitates the precise encoding of an animal’s environment and location.
@@
Derivation of Grid Cell Positional Encoding
Inner Product Translational Invariance
@@
$\sum_{g\in\mathcal{G}}\left\langle g(\boldsymbol{x}),g(\boldsymbol{y})\right\rangle_{\mathcal{H}}$ $=\sum_{i}|c_{i}|^{2}e^{j\boldsymbol{k}_{i}^{T}\boldsymbol{d}}$ (7)
@@
Note this function is also a translational invariant kernel for $\boldsymbol{x}$ and $\boldsymbol{y}$. Simultaneuously, in its form it scale to the inverse Fourier transformation of the probability distribution of wave vectors. According to Bochner theorem (Bochner 2005), we know this kernel, as the inverse Fourier transformation of a non-nagative measurement, is positive definite. Therefore, we can apply random Fourier feature (RFF) (Rahimi and Recht 2007) technique on this shift-invariant kernel, which demands positive definite kernel as the prerequisite. We denote the $\kappa(\boldsymbol{x},\boldsymbol{y})$ as this kernel and we can know:
@@
$=\frac{1}{\sqrt{D}}\left(\begin{array}[]{c}e^{j\boldsymbol{k}_{1}^{T}\boldsymbol{x}}\\ e^{j\boldsymbol{k}_{2}^{T}\boldsymbol{x}}\\ \vdots\\ e^{j\boldsymbol{k}_{D}^{T}\boldsymbol{x}}\end{array}\right)^{T}$
$\quad\cdot\begin{bmatrix}|c_{1}|^{2}&&&\\ &|c_{2}|^{2}&&\\ &&\ddots&\\ &&&|c_{D}|^{2}\end{bmatrix}$
$\quad\cdot\left(\begin{array}[]{c}e^{-j\boldsymbol{k}_{1}^{T}\boldsymbol{y}}\\ e^{-j\boldsymbol{k}_{2}^{T}\boldsymbol{y}}\\ \vdots\\ e^{-j\boldsymbol{k}_{D}^{T}\boldsymbol{y}}\end{array}\right)\frac{1}{\sqrt{D}}$
@@
We propose that the brain represents Euclidean space using multi-dimensional Fourier analysis, akin to the auditory system’s processing of sound waves and the visual system’s processing of light waves. By integrating oscillations of various spatial frequencies, the grid cell encoding system can efficiently represent spatial information with high resolution. Furthermore, the Random Fourier Features (RFF) approach imposes no constraints on the dimensionality of locations $\boldsymbol{x}$ and $\boldsymbol{y}$. This suggests that the Fourier representation of grid cells can encode location information in arbitrarily high-dimensional spaces.
Optimal Ratio of Grid Scale under Economy Principle
In the derivation above, we assumed the wave vector (i.e., the preferred direction) follows a probability distribution, such as the uniform distribution often used in RFF. However, neuroscience evidence shows that the medial entorhinal cortex consists of several modules. Within the same module, the direction and period scale of the cell grids remain consistent, but these discretized scales shrink along the ventral-dorsal axis (Stensola et al. 2012), representing a transition from coarse, low-frequency to granular, high-frequency spatial information. The periodic ratio of the adjacent scale remains around 1.5, as estimated by both animal measurements and computational simulations (Stensola et al. 2012; Banino et al. 2018). This hierarchical spatial representation demands significantly fewer grid cells while maintaining fine spatial resolution, or say the economy principle (Wei, Prentice, and Balasubramanian 2015).
The challenge now is to determine the optimal ratio of the grid scale, which is crucial for constructing a positional encoding system. Recall that in the original positional encoding of the Transformer architecture (Vaswani et al. 2017), the encoding spatial frequency is calculated as $\omega_{k}=\frac{1}{10000^{2k/d}}$, forming a geometric series of wavelengths. The rationale behind this choice remains unclear, and we anticipate that computational theories of grid cells can provide insights into frequency determination.
@@
Here we provide a generalized derivation based on the proof in (Wei, Prentice, and Balasubramanian 2015). We aim to represent p-dimensional Euclidean space with the least number of grid cells, reflecting the principle of economy. Assume there are $m$ scales (or layers) of the grid cells. The period of the $i$-th module layer ($i=1,2,\ldots,m$), or the distance between the adjacent grid vertices, is denoted as $\lambda_{i}$. The period increase as $i$ becomes larger. The scale ratio of the adjacent layers is $r=\lambda_{i+1}/\lambda_{i}>1$. A grid field is a hyperspherical area centered around a grid vertex in the space, with its diameter denoted as $l_{i}$. For 2D space, a grid field is a circular-plate area activating the grid cell. Clearly, to cover the whole space without omissions, more than one grid cell with different phases are required. The larger the ratio of $\lambda_{i}$ to $l_{i}$, the more grid cells are needed. If there must be at least $d$ cells responding to a spatial location, the total number of grid cells across all layers can be expressed as $N$:
@@
In the original Transformer architecture, the adjacent scale of positional encoding is described by Equation 12. For instance, when the encoding dimension $d_{\text{model}}$ is 128, the difference in adjacent scales is $10000^{\frac{2}{128}}\approx 1.457$, whereas for an encoding dimension of 512, the difference in adjacent scales is $10000^{\frac{2}{512}}\approx 1.036$. This relationship can be expressed as follows:
$\rho_{\text{ratio}}=\frac{10000^{\frac{2(i+1)}{d_{\text{model}}}}}{10000^{\frac{2i}{d_{\text{model}}}}}=10000^{\frac{2}{d_{\text{model}}}}$ (12)
As discussed, to avoid ambiguity in positional encoding, the maximum scale ratio for encoding in an $n$-dimensional space is $r=\sqrt[p]{e}$. Consequently, the minimum encoding dimension can be determined when the base is 10000. To achieve an adjacent scale difference of $p]{e}$, we can establish the following inequality:
$10000^{\frac{2}{d_{\text{model}}}}\leq\sqrt[p]{e}$ (13)
Taking the logarithm of both sides and solving for $d_{\text{model}}$:
$\frac{2}{d_{\text{model}}}\log(10000)\leq\frac{1}{p}\log(e)$ (14)
$d_{\text{model}}\geq\frac{2\log(10000)p}{\log(e)}$ (15)
This allows us to calculate the minimum encoding dimension $d_{\text{model}}$ that satisfies the specified adjacent scale difference.
Proposal of Grid-cell Positional Encoding (GridPE)
After demonstrating the capability of representing relative movement in space through inner product calculations and determining the optimal ratio of encoding periods, we introduce a unified positional encoding scheme for attention-based architectures. Inspired by neuroscientific principles of grid cells, our method extends the foundational ideas presented by (Su et al. 2024). While Su’s model combines absolute and relative positional encodings within a fixed framework, our schema advances this by incorporating grid cell-inspired encodings that inherently support extrapolation to higher dimensions and maintain translational invariance. Notably, in one-dimensional space, where GridPE can be expressed using a single Fourier basis, our encoding scheme is equivalent to rotational positional encoding. This demonstrates the flexibility of our approach and its alignment with established encoding strategies, further enhancing its applicability across various architectural contexts. According to Su’s model, to accommodate both absolute and relative positional encoding, the inner product of the query $\boldsymbol{q}$ and key $\boldsymbol{k}$, which represents their similarity, can be modified as follows:
@@
$\boldsymbol{W}_{m}=\left[\begin{array}[]{c}\sum_{i}c_{i}e^{j\boldsymbol{\omega}_{i,1}^{T}m}\\ \sum_{i}c_{i}e^{j\boldsymbol{\omega}_{i,2}^{T}m}\\ \vdots\\ \sum_{i}c_{i}e^{j\boldsymbol{\omega}_{i,d}^{T}m}\end{array}\right]$ (17)
@@
$\boldsymbol{R}_{m}=\begin{bmatrix}\boldsymbol{A}_{1}&\boldsymbol{0}&\dots&\boldsymbol{0}\\ \boldsymbol{0}&\boldsymbol{A}_{2}&\dots&\boldsymbol{0}\\ \vdots&\vdots&\ddots&\vdots\\ \boldsymbol{0}&\boldsymbol{0}&\dots&\boldsymbol{A}_{d/2}\\ \end{bmatrix}$ (18)
$\boldsymbol{A}_{k}=\begin{bmatrix}\Re\left(e^{j\boldsymbol{\omega}_{k,1}^{T}m}\right)&-\Im\left(e^{j\boldsymbol{\omega}_{k,1}^{T}m}\right)\\ \Im\left(e^{j\boldsymbol{\omega}_{k,1}^{T}m}\right)&\Re\left(e^{j\boldsymbol{\omega}_{k,1}^{T}m}\right)\\ \end{bmatrix}$ (19)
@@
We can directly use the real part of the inner product between two sets of ”complex” position encoding as the encoding product.
@@
Experiments
Property of Distance Decay
We first display that our positional encoding approach exhibits the property of distance decay in any high-dimensional space. This implies that the association between two positions diminishes as the distance between them increases. We select a square area in two-dimensional space, calculating the positional encoding vector within the area as shown in Equation 17, and compute the inner products between the central ”core” point and uniformly sampled points. Given that we have proven the translational invariance of the inner product, our findings remain valid when any other point is chosen as the core. Figure 3 presents the inner product surface plot and a profile plot with $x$ fixed at $0.5$.